OpenStudy (anonymous):

How do you graph y=x^2-4x+3 I need the vertex, axis of symmetry, y and x axis as well will give medal

OpenStudy (anonymous):

complete the square on R.H.S.

OpenStudy (anonymous):

\[y=x^2-4x+4-4+3=(x-2)^2-1\] \[(x-2)^2=y+1\] it is an upward parabola with vertex (2,-1)

OpenStudy (anonymous):

axis of symetry is x-2=0 or x=2

OpenStudy (anonymous):

it cuts y-axis at (0,3)

OpenStudy (anonymous):

it cuts x-axis where y=0 \[(x-2)^2=1,x-2=\pm1,x=?\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw